OpenStudy (roseliberty):

i'll fan and medal Finding derivatives??? Using the rules of derivatives, find derivative of this function: f(x)=(6x^2+1)/(5x^2-2)?

myininaya (myininaya):

Have you tried using the quotient rule?

myininaya (myininaya):

let u=6x^2+1 and v=5x^2-2 can you find both u' and v' ?

myininaya (myininaya):

\[(\frac{u}{v})'=\frac{u' v-v' u}{v^2}\]